OpenStudy (anonymous):

Ms. Smith invested $13,000 in two accounts, one yielding 4% interest and the other yielding 9%. If she received a total of $770 in interest at the end of the year, how much did she invest in each account?

jimthompson5910 (jim_thompson5910):

Hint: Solve the system x+y = 13000 0.04x + 0.09y = 770

OpenStudy (anonymous):

What amount was invested a t 4%?

jimthompson5910 (jim_thompson5910):

to find this answer, solve for x

jimthompson5910 (jim_thompson5910):

x+y = 13000 y = 13000 - x --------------------- 0.04x + 0.09y = 770 0.04x + 0.09(13000 - x) = 770 keep going to solve for x

OpenStudy (anonymous):

@chicashley : Ms. Smith invested a total of 13000 dollars in two accounts. This means that she invested a certain portion of 13000 in the first account(call it account in Bank A) and the remaining in the second account.(account in Bank B). This means that Amount in Bank A plus Amount in Bank B equals 13000. This is what @jim_thompson5910 meant by x+y = 13000 Next, Bank A gives an interest of 4% and Bank B gives an interest of 7%. The total amount of money Ms Smith received in interest was 770 dollars. This means that 4 percent of the amount in Bank A plus 7 percent of the amount in Bank B is equal to 770 dollars, or, 4%*Amount in Bank A + 7%*amount in Bank B = 770 This is what @jim_thompson5910 meant by 0.04x + 0.07y = 770. Now, you have two equations: x+y = 13000 and 0.04x+0.07y=770. solve these two equations to get x and y. And keep in mind what x and y represent. x represents the amount of money in Bank A y represents the amount of money in Bank B

OpenStudy (anonymous):

$4480. @ 4% 8520. @ 9%